OpenStudy (anonymous):

Solve please. 1. Nora’s leaf collection consisted of 150 leaves in August. In December, her collection grew to 177 leaves. Find the percent increase between August and December. (Points : 1) 9% 12% 15% 18%

OpenStudy (sleepyjess):

150/100=1.5 so 1.5 leaves is equal to 1% 177-150=27 her collection grew by 27 leaves 27/1.5=percent

OpenStudy (luckycoins888):

15% i think

OpenStudy (sleepyjess):

how did you get to 15%? @luckycoins888

OpenStudy (anonymous):

more help 2. The cost per minute for a conference phone call, originally $2.60, increased by 30%. What is the new price? (Points : 1) $2.90 $3.20 $3.38 $3.48 Question 3.3. Selena buys several shares of stock at $32 per share and sells them at $28.32 per share. What is the percent of decrease in the price of the stock? (Points : 1) 12.0% 11.5% 11.0% 10.5% Question 4.4. Katie earned $1890 in 7 years on an investment at a 6% simple interest rate. How much was Katie’s investment? (Points : 1) $945 $793.80 $4500 $9000 Question 5.5. Find the annual interest rate if a $5000 investment earns $550 annual simple interest in one year. (Points : 1) 10.5% 11% 11.5% 12% thank you

OpenStudy (sleepyjess):

is this a test?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

help

OpenStudy (sleepyjess):

im sorry but it is against the CoC to help on tests or ask questions from tests \(\large\bf\color{blue}{If~you~are~unsure~of~the~rules~please~review~the~\\Code~of~Conduct~}\href{http:///openstudy.com/code-of-conduct}{\large\bf\color{red}{here}}\)

OpenStudy (sleepyjess):

how does .18=.12?

OpenStudy (michele_laino):

@sleepyjess are you saing to me? if so i'm going to delete my answer, tell me, please

OpenStudy (sleepyjess):

yes

OpenStudy (michele_laino):

Ok!, done!, thanks @sleepyjess